OpenStudy (anonymous):

In the triangle below, b = _____. If necessary, round your answer to two decimal places.

OpenStudy (ibbutibbu.):

The triangle might me helpful. And the link you posted doesnt work.

OpenStudy (ibbutibbu.):

For this we can use the angle sum rule, which states that the sum of all angles in a triangle is 180 degrees. So, 46.7 + 48 is 94.7. Now, you can subtract that from 180. And that is your answer. Can you tell me what that is?

OpenStudy (anonymous):

so b =85.3

OpenStudy (ibbutibbu.):

YES!!! Great job!!!

OpenStudy (ibbutibbu.):

Do you understand the concept now?

OpenStudy (anonymous):

sort of

OpenStudy (aum):

Use the Law of Sines: b / sin(B) = a / sin(A) b = ( a * sin(B) ) / sin(A) ----- (1) a = 44 A = 46.7 degrees B = 180 - (A+C) = 180 - (46.7 + 48) = ? Plug the values into (1) and compute b

OpenStudy (ibbutibbu.):

@aum but am i correct? Becuase i have not learned sine and cosine yet :|

OpenStudy (aum):

You have found upper case B which represent the angle B in the triangle. And B = 85.3 degrees is correct. Here the question asks for lower case 'b' which represents the length of the side opposite to the angle B.

OpenStudy (aum):

So the question here asks for the length of the side AC.

OpenStudy (ibbutibbu.):

No, it asks for angle B

OpenStudy (aum):

"In the triangle below, b = _____. If necessary, round your answer to two decimal places." It asks for lower case 'b'. Look in the diagram what lower case 'b' represents. It is the length of the side AC.

OpenStudy (aum):

Refer to my first reply: b = ( a * sin(B) ) / sin(A) ----- (1) b = ( 44 * sin(85.3) ) / sin(46.7) = ? (set your calculator to degree mode before calculating above)

OpenStudy (ibbutibbu.):

Oh ok then

OpenStudy (anonymous):

hey ibb you arent teaching anything when some one asks a question fro apex they want an answer not help freak

OpenStudy (anonymous):

@TrippleGRIN777 I believe answering with no teaching is in direct violation of the rules. OpenStudy is a tutoring tool, not for you to cheat off of.